# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

### v1

@@ 4 Discussion
Stand-dependent dispersal, an Allee effect, and time-scale separation are jointly necessary for realistic mountain pine beetle outbreaks. Time-scale separation (i.e., trees grow slower than beetles) ensures that killed trees cannot be immediately replaced by new mature pines, thus preventing equilibrium dynamics. Stand-dependent dispersal (i.e., beetles disperse farther through sparse forests) leads to high levels of MPB mortality, and the Allee effect finishes the MPB population off, bringing it to the endemic state. These two inhibitory processes (Stand-dependent dispersal and the Allee effect) explain why MPB populations spend the vast majority of time in their endemic state, and why subsequent outbreaks are so explosive: with MPB populations suppressed, large trees accumulate, providing proverbial fuel for the beetle fire.
The process of stand-dependent dispersal explains why beetles don’t overexploit trees. It is perhaps trivial to say that if beetles disperse away and die before they have killed all the trees, then the beetles won’t kill all the trees. However, the lack of over-exploitation, or equivalently, the seemingly premature “bust” phase of the “boom-bust cycle”, is a ubiquitous feature of outbreaks (Fig. 5) and is thus worth connecting to MPB life-history. This also suggests an optimistic perspective: outbreaks are much less destructive than they theoretically could be. Stand-dependent dispersal probably occurs because chemo-sensory constraints prevent beetles from aggregating on small trees (Section 1.2.2), even though beetles have positive fitness when they do successfully attack small trees. Without stand-dependent dispersal, the system would reach a stochastic equilibrium, creating a forest with few large pine trees.
Classic predator-prey theory predicts that long-lived prey should make predator-prey cycles unlikely. One might naively expect that the time-scale separation between beetles and trees would allow MPB to grow rapidly enough to prevent an unchecked proliferation in pine abundance, thus preventing the prey-escape phase of a population cycle (De Roos et al.,, 1990) wherein the prey species can grow unimpeded by the predator. Even though pines live much longer than beetles, we still see population cycles in the counterfactual simulations (when the Allee effect is turned off). This can be explained by the phenomenon of cohort resonance (Bjørnstad et al.,, 2004). During an outbreak, many pines are killed at roughly the same time. Subsequently, pines are recruited at the same time, and then become susceptible to MPB at the same time. Thus, there are a brief few years where the pines that MPB can “see” are growing faster than exponentially.
@@
Our model accurately predicts the return time of MPB outbreaks, which is somewhere between 20–50 years according to tree-ring-based reconstructions. Our counterfactual simulations with no Allee effect, resulting in classic population cycles, predict an outbreak period that is too short (approximately 20 years). Other models in the literature claim an outbreak periodicity that is too long (Powell and Bentz,, 2009; Duncan et al.,, 2015; Brush and Lewis,, 2023).
@@
The period of MPB outbreaks is primarily determined by the time it takes lodgepole pine to grow into a susceptible age class. Thus, if MPB were to kill all of the trees, then outbreaks would occur approximately every 55 years: 50 years for the trees to become susceptible and several more years for the outbreak and subsequent recruitment of pine saplings. However, there are two real-world complications. First, stand-dependent dispersal results in a large number of medium-sized trees surviving the outbreak, so the observed periodicity of 35 years partially reflects the time it takes for medium-sized trees to become large trees, e.g., it takes approximately 25 years for an 8-inch tree to become a 12-inch tree. The 2000s outbreak in British Columbia killed more medium-sized trees than usual, suggesting that it will take longer than usual for the next outbreak to occur. Second, large trees are a pre-requisite for outbreaks, but widespread outbreaks require widespread perturbations that allow several populations to synchronously break out of the low-density endemic state. This perturbation is usually realized as drought conditions. Tree growth and climate patterns lead to a right-skewed distribution of waiting times between outbreaks, where the minimum time is determined by tree growth, and the shape of the long right tail is determined by drought probability.
The life-history processes of size-dependent fecundity and size-dependent attack preference do not explain the qualitative behavior of MPB dynamics, but they do explain about 25% of the magnitude of outbreaks (i.e., the peak number of beetles). Size-dependent attack preference focuses the beetles on the large trees. Under size-dependent fecundity, the large trees are the most productive (from the beetles’ perspective), so more beetles are produced before most of the susceptible trees are eaten up, triggering stand-dependent dispersal and the subsequent “bust” phase of the outbreak.
@@
The peak number of beetles is a highly relevant statistic, since it is likely the main driver of MPB range expansion. Only a minuscule fraction of emerging beetles undergo long-distance dispersal (estimated at 0–2%), so outbreak expansion via long-distance dispersal requires severe local infestations, such that emigrating beetles are numerous enough to mass attack trees — and in doing so, overcome the Allee effect — at their destination. In other words, range expansion requires that peak beetle density exceeds the donor threshold (terminology introduced by Johnson et al.,, 2006). The concept of a donor threshold may explain why outbreaks are spatially correlated at greater distances in British Columbia than in the western United States (U.S. pine forests generally have less pine biomass $\rightarrow$ less peak beetles), and why MPB spread into central Alberta via long-distance dispersal from British Columbia, rather than slowly spreading up the east side of the Rocky Mountains.
Through our model-building process and our simulation study, we have gathered insights about the determinants of MPB’s distinctive population dynamics. Stand-dependent dispersal and the Allee effect are essential, whereas size-dependent attack preference and size-dependent fecundity are less important. Beetle attack density may increase throughout an outbreak, but this has a negligible effect on population dynamics (Appendix B.2). MPB’s preference for large-diameter trees may become exacerbated when there are many beetles, but this has a negligible effect on population dynamics, at least in the epidemic phase of an outbreak (Appendix B.1).
@@
Our research offers a revised perspective on the density-dependent dynamics of MPB, challenging established assumptions. In previous work, the number of new infestations is assumed to be proportional to both beetles and trees (e.g. Burnell,, 1977; Heavilin et al.,, 2007; Křivan et al.,, 2016), as in the classic Lotka-Volterra model. Yet, our analysis reveals a different perspective for cases of intermediate-to-high tree density (i.e., before stand-dependent dispersal kicks in): the number of new infestations is only proportional to beetles, reflecting the fact that outbreaking beetles are limited by their ability to overcome tree defenses, not by their ability to find susceptible trees. While overcompensatory density dependence has been a common assumption in past studies (e.g. Goodsman et al.,, 2017, 2018; Lewis et al.,, 2010; Duncan et al.,, 2015), we argue that MPB experiences near-perfect compensatory density dependence (see Section 1.2.2). Previous research has assumed that the probability of a successful mass attack depends on the quotient of beetle and tree density (e.g., Burnell,, 1977; Goodsman et al.,, 2016; Brush and Lewis,, 2023). Instead, we propose that beetle density alone is a more significant determinant of the success probability, as detailed in Appendix B.3.
@@
Putting all of these results together, we can simplify our simulation model (Section 2.3) to formulate a minimal mechanistic model of MPB dynamics. Because size-dependent preference and size-dependent fecundity are non-essential, we drop explicit age structure and only track the density of all susceptible trees (i.e., trees with DBH $>$ 8 inches). All susceptible trees are equally preferred for mass attacks, and all killed trees give rise to the same number of beetles. For simplicity, stochasticity is eliminated. The minimal mechanistic model can be written as
@@
$\underbrace{S_{t+1}}_{\mathclap{\text{Susceptible trees per acre}}}=\overbrace{g(K^{*}-S_{y})}^{\mathclap{\text{Recruitment of susceptible trees}}}-\underbrace{A_{y}\times(1-e^{\frac{-2.3A_{y}}{a}})\times\overbrace{\gamma^{*}}^{\mathclap{\text{Trees killed per succ. beetle attack}}}}_{\mathclap{\text{Trees killed by MPB}}}.$ (18)
@@
Here, $A_{y}$ is the female attack density in the $y^{\text{th}}$ year, $S_{y}$ is the density of susceptible trees, $S_{y}^{\prime}$ is the density of susceptible trees remaining after beetle attack (i.e., $S_{y}^{\prime}=S_{y}-A_{y}\times(1-e^{\frac{-2.3A_{y}}{a}})$), the parameter $\zeta^{*}$ is the average emerging females per killed tree, $\gamma^{*}$ is the average number of killed trees per successful attack, $K^{*}$ is redefined as the carrying capacity of the susceptible trees, $\epsilon$ is the now-explicit number of endemic female beetles, $g$ modulates the rate of recruitment into the class of susceptible trees, and all other parameters have their original meanings (see Table A.1).
@@
The minimal model has several purposes. 1) It is descriptive, compactly representing the density-dependent dynamics of MPB and lodgepole pine. 2) It demonstrates that there are two stable equilibria: an endemic equilibrium with few beetles, and an epidemic equilibrium with many beetles. This bistability is typical of systems with a strong Allee effect (Taylor,, 2012), so it is no surprise that the endemic equilibrium disappears when the Allee effect is turned off (Fig. 20C) 3) It shows the MPB/lodgepole system is an instance of an excitable dynamical system, where perturbations can lead to long excursions in phase space, before returning to a resting state (Fig. 20A). Typically, excitable systems have a single stable equilibrium (Strogatz,, 1994, pg. 118), whereas our system has two. However, the endemic equilibrium is globally-attracting as a practical matter, since the epidemic equilibrium has a minuscule basin of attraction (Fig. D.6).
@@
Figure 20: Phase portraits of the minimal mechanistic model (Eq. 17 & Eq. 18), for scenarios where stand-dependent dispersal and the Allee effect have been turned on or off. Most model parameters were set to their respective posterior means (see Table 2). The exceptions are $\gamma^{*}=1/360$ and $\zeta^{*}=1300$, which are approximately the respective means (across trees) of attacks and emergences in the Klein et al., dataset. The carrying capacity’s value, $K^{*}=130$, is approximately the density of trees with $\text{DBH}\geq 8\text{in}$ in the Klein et al., dataset. The recruitment parameter’s value, $g=0.036$, was obtained by solving the differential equation $dS(t)/dt=g(K-S(t))$ with initial condition $S(0)=0$, then solving for the $g$ where 90% of carrying capacity is attained in $t=50$ years; lodgepole pine grows attains an 8-inch DBH and becomes susceptible after approximately 50 years.
While our models thoroughly describe the density-dependent dynamics of MPB, they do not account for the impacts of environmental heterogeneity or stratified dispersal. There are additional un-modelled processes, including temperature-dependent phenology, beetle vigor depending on natal tree size, competition with other bark beetle species, and predation by natural enemies like woodpeckers and parasitoid wasps. While we are confident in the structure of our model, parameter estimates primarily rely on data from a single outbreak at a single geographic location. Thus, our models should not be used to make precise quantitative predictions.
@@
Through the development and analysis of our age-structured simulation model, we have shed light on the life-history traits underlying outbreaks, and challenged established ideas in the MPB modelling literature. Our findings are succinctly summarized in a minimal mechanistic model, which incidentally establishes the MPB/lodgepole system as one of the first empirical examples of an excitable dynamical system in ecology. This work underscores the importance of detailed case studies and looking beyond traditional predator-prey models to understand complex population dynamics. Future research should leverage our description of dynamics to model the extensive Aerial Overview Survey data from the United States and British Columbia, both to forecast outbreaks and to understand large-scale dispersal patterns.

### v2

@@ 4 Discussion
Excitable population dynamics are related to five features of bark beetle ecology: stand-dependent dispersal mortality, an Allee effect, time-scale separation between beetle and tree life cycles, tree size-dependent fecundity, and a large-tree preference. Time-scale separation (i.e., trees grow slower than beetles) ensures that killed trees cannot be immediately replaced by new mature pines, thus preventing equilibrium dynamics. Stand-dependent dispersal (i.e., beetles disperse farther through sparse forests) leads to high levels of mortality, which is exacerbated by the Allee effect (i.e., not enough beetles to mass attack healthy trees), bringing beetle populations to an endemic, low-density state. The two inhibitory processes — stand-dependent dispersal mortality and a strong Alle effect — explain why populations spend the vast majority of time in their endemic state, and why subsequent outbreaks are so explosive: with suppressed beetle populations, large trees accumulate, providing proverbial fuel for the beetle fire.
The two remaining traits, tree size-dependent fecundity and a large-tree preference, explain the large magnitude of outbreaks, i.e., the peak number of beetles. Larger trees provide more resources for beetle reproduction and are preferentially attacked, leading to rapid population growth during outbreaks. While not essential for MPB’s excitable population dynamics due to MPB’s high baseline fecundity, these traits may be crucial for other bark beetle species that lie on the boundary between pulse-driven and irruptive behaviors. Bark beetles can be categorized as either “outbreaking” or “non-outbreaking”, with the former category further divided into “pulse-driven” and “irruptive” (equivalent to excitable) categories.
4.1 Bark beetle comparative analysis
Aggregation behavior is a key distinguishing factor between outbreaking and non-outbreaking bark beetle species. Non-outbreaking species are poor aggregators, with few individuals attacking simultaneously and pheromones used primarily for mate-finding. Examples include the greater European spruce beetle (Dendroctonus micans) and the black turpentine beetle (Dendroctonus terebrans); more are listed in Reeve et al., (2012) Table 2. To survive in living trees, non-outbreaking beetles often have adaptations such as resistance to toxic tree resin (Weed et al.,, 2015). Interestingly, Dendroctonus micans adults do not aggregate, but larvae do (Grégoire,, 1988; Deneubourg et al.,, 1990), leading to greater larval survival in small pockets of tree phloem (Storer et al.,, 1997). Although most bark beetles are non-outbreaking (Wood,, 1982), aggregation and mass-attacks appear to be the ancestral state for Dendroctonus bark beetles (Reeve et al.,, 2012).
The distinction between pulse-driven and irruptive species is less clear-cut. Instead, differences lie in the magnitude of various life history traits and environmental factors. For example, the Douglas-fir beetle (Dendroctonus pseudotsugae), a pulse-driven species, prefers large trees that are downed (Hood and Bentz,, 2007) or recently defoliated by the Douglas-fir tussock moth, Orgyia pseudotsugata (Wright et al.,, 1984; Negrón et al.,, 2014). During outbreaks, it may attack healthy trees, but these are typically adjacent injured trees (Cunningham et al.,, 2005; Cole et al.,, 2022). This behavior, contrasting with the irruptive MPB’s strong preference for healthy trees, contributes to the Douglas-fir beetle’s classification as a pulse-driven species.
The western balsam bark beetle (WBBB; Dryocoetes confusus) presents an intriguing case that further blurs the lines between these categories. The WBBB likely has excitable population dynamics: it is generally considered irruptive (e.g. Lantschner and Corley,, 2023) and causes large destruction in response to drought perturbations (Maclauchlan et al.,, 2023). Further, WBBB shares many life history traits with the MPB, including a preference for large trees, tree-size-dependent fecundity, and the mass-attack-based Allee effect. Despite these similarities, their population dynamics look quite different. WBBB often attacks less than 5% of a given stand in one year, with chronic infestations slowly killing many mature trees over time (Maclauchlan et al.,, 2003; McMillin et al.,, 2003). One possible explanation is that WBBB carries a fungus, Grosmannia dryocoetidis, that is truly pathogenic to host trees (Molnar,, 1965; Zipfel et al.,, 2006). Therefore, sub-critical or initially unsuccessful attacks can kill trees simply by inoculating the fungus. Another possible explanation is that WBBB’s primary host, subalpine fir, often grows in mixed-age stands (Antos and Parish,, 2002; Maclauchlan,, 2016) wherein younger trees are less susceptible. This contrasts with MPB’s primary host, lodgepole pine, which typically experiences clearcut logging or stand-leveling fires, resulting in even-aged stands prone to explosive outbreaks. The highly clustered nature of WBBB’s outbreaks (Howe et al., 2022a, ) suggest it may lack stand-dependent dispersal mortality.
A beetle species’ status as irruptive or pulse-driven can change over space and time in response to environmental conditions. One example is the red turpentine beetle (Dendroctonus valens), which is considered non-outbreaking in its native range in North America but has become irruptive in China. This shift is attributed to novel fungal symbionts aiding host tree mortality, the higher susceptibility of the Chinese host Pinus tabuliformis, more attractive host volatiles, prevalent monoculture plantations, and warmer weather (Yan et al.,, 2005; Sun et al.,, 2013). This case highlights the importance of considering traits (e.g., aggregation and large-tree preference due to attractive host volatiles), abiotic factors (e.g., climate), and idiosyncratic elements (e.g., fungal symbionts) in determining outbreak potential.
The European spruce beetle (Ips typographus) provides another interesting comparison to MPB. I. typographus experiences significantly higher intraspecific competition among larvae, resulting in consistently lower fitness (KärvEmo and Schroeder,, 2010). This difference may explain why the European spruce beetle is less destructive than MPB (KärvEmo and Schroeder,, 2010), despite both species being considered irruptive. The likely mechanism behind this disparity is I. typographus’ much lower emission of the anti-aggregation pheromone verbenone (Birgersson et al.,, 1984; Ramakrishnan et al.,, 2022). Another way to conceptualize the disparity (in terms of our “biological features”), is that the European spruce beetle lacks strong tree size-dependent fecundity, since high competition undermines the benefits of thicker phloem.
Environmental conditions and interspecific interactions significantly influence outbreak potential. The spruce beetle (Dendroctonus rufipennis) in Alaska’s interior rarely outbreaks despite favorable conditions. Less precipitation in the interior leads to lower phloem moisture, thus favoring competitive non-eruptive insects like Ips perturbatus (Werner et al.,, 2006). Similarly, the roundheaded pine beetle (Dendroctonus adjunctus) outbreaks in areas where trees are stressed by drought and other bark beetles, specifically western pine beetles (Dendroctonus brevicomis) and Ips species (Negron et al.,, 2000; Negrón,, 2008).
We conducted a comparative analysis of bark beetle damage using aerial survey data from western North America and maps of host-tree biomass (Appendix D). Our analysis reveals that classification as ”irruptive” is a necessary condition for causing extensive damage, but the total amount of beetle damage scales linearly with the total biomass of host trees (Fig. 9). Previous research has found an approximately linear relationship between tree biomass and tree damage across sites, within species (Negron et al.,, 1999; Negron and Klutsch,, 2017); here we show that this relationship also applies between species. Together, both life history traits and environmental factors determine if a bark beetle population is capable of overcoming an Allee effect (i.e., overcoming the defenses of healthy trees) and growing explosively; then landscape-level factors determine the overall impact of bark beetle outbreaks. This analysis adds a layer of nuance to Rudinsky’s 1962 claim that the “Abundance of suitable breeding material has recently been definitely identified as the decisive factor in outbreaks of bark beetles.”
Figure 9: Total bark beetle damage is determined by life-history traits and landscape characteristics. Cumulative damaged area is a linear function of the total biomass of host trees, across British Columbia and the western United States (i.e., regions 1–6 as defined by the US Forestry Service); however, this relationship only holds true for bark beetles classified as irruptive. Methodological details in Appendix D.
The western pine beetle (Dendroctonus brevicomis) provides an illustrative example of how landscape characteristics can influence the impact of even highly irruptive species. Unlike MPB, which can attack any pine species within its native range, the western pine beetle’s outbreaks are limited due to its high specificity to ponderosa pine (Homicz et al.,, 2022). Recent trends in southern pine beetle (Dendroctonus frontalis) damage illustrate how management practices can affect landscape characteristics and total beetle impact. Despite being considered irruptive, the total area affected by southern pine beetle has declined significantly over time (Fettig et al.,, 2022). Asaro et al., (2017) attribute this decline to silvicultural practices, namely rapidly removing infested trees and limiting connectivity between high-risk host stands.
Other bark beetle species lack the unique life history of MPB, either in kind or in degree, resulting in pulse-driven outbreaks or less destructive irruptive outbreaks. We have examined examples of bark beetles lacking MPB’s key traits: the great European spruce beetle lacks the ability to aggregate (implicating the Allee effect), the western balsam bark beetle possibly lacks stand-dependent dispersal mortality, and the European spruce beetle lacks size-dependent fecundity. The Douglas-fir beetle prefers large trees Sturdevant et al., (2022), but prefers these to be downed or otherwise unhealthy. Consequently, these species have slow-moving outbreaks (e.g., the western balsam bark beetles) or are unable to achieve self-driving growth without sustained perturbations.
4.2 Mountain pine beetle dynamics
@@
Our model accurately predicts the return time of MPB outbreaks, which is somewhere between 20–50 years according to tree-ring-based reconstructions. Our counterfactual simulations with no Allee effect, resulting in classic population cycles, predict an outbreak return time that is too short (approximately 20 years). Other models in the literature claim an outbreak return time that is too long (Powell and Bentz,, 2009; Duncan et al.,, 2015; Brush and Lewis,, 2023).
@@
The period of MPB outbreaks is primarily determined by the time it takes lodgepole pine to grow into a susceptible age class. In an idealized scenario where MPB kills all trees, outbreaks would occur approximately every 55 years: 50 years for new trees to become vulnerable, plus a few years for the outbreak and subsequent growth of pine seedlings. However, there are two real-world complications. First, stand-dependent dispersal mortality results in a large number of medium-sized trees surviving the outbreak, so the observed periodicity of 35 years partially reflects the time it takes for medium-sized trees to become large trees, e.g., it takes approximately 25 years for an 8-inch tree to become a 12-inch tree. The 2000s outbreak in British Columbia killed more medium-sized trees than usual, which suggests a longer-than-usual wait time for the next outbreak. Second, widespread outbreaks require widespread perturbations (usually drought), which are functionally stochastic. Thus, tree growth and climate patterns imply a right-skewed distribution of waiting times between outbreaks, where the minimum time is determined by tree growth, and the shape of the right tail is determined by drought probability.
@@
Tree size influences both MPB fecundity and attack preference, explaining about 40% of peak beetle numbers. The outbreak peak is relevant because it is thought to drive MPB range expansion (Bleiker,, 2019). Only a minuscule fraction of emerging beetles undergo long-distance dispersal (estimated at 2.5%; Safranyik et al.,, 1992), so outbreak expansion via long-distance dispersal requires severe local infestations, such that emigrating beetles are numerous enough to mass-attack trees at their destination. In other words, range expansion requires that peak beetle density exceeds the donor threshold (terminology introduced by Johnson et al.,, 2006).
The donor threshold concept may explain regional patterns of MPB spread. Outbreaks are spatially correlated at greater distances in British Columbia than in the western United States, where lower pine biomass leads to smaller peak beetle populations and fewer successful long-distance dispersers. The donor threshold may also explain why MPBs expanded into central Alberta from British Columbia, rather than gradually moving northward from Montana along the eastern Canadian Rockies. British Columbia is more climatically suitable than the eastern Rockies (Safranyik et al.,, 2010), thus leading to large MPB populations and successful long-distance dispersal.
@@
Our research offers a revised perspective on the density-dependent dynamics of MPB, challenging established assumptions. In previous work, the number of new infestations is assumed to be proportional to both beetles and trees (e.g. Burnell,, 1977; Heavilin et al.,, 2007; Křivan et al.,, 2016), as in the classic Lotka-Volterra model. Yet, our analysis shows that for intermediate-to-high tree density (i.e., before stand-dependent dispersal mortality kicks in) the number of new infestations is only proportional to beetles, reflecting the fact that outbreaking beetles are limited by their ability to overcome tree defenses, not by their ability to find susceptible trees. While overcompensatory density dependence has been a common assumption in past studies (e.g. Goodsman et al.,, 2017, 2018; Lewis et al.,, 2010; Duncan et al.,, 2015), we argue that MPB experiences near-perfect compensatory density dependence (see Section 1.2.2). Previous research has assumed that the probability of a successful mass attack depends on the quotient of beetle and tree density (e.g., Burnell,, 1977; Goodsman et al.,, 2016; Brush and Lewis,, 2023). Instead, we propose that beetle density alone is a more significant determinant of the success probability, as detailed in Appendix B.5.
@@
Putting all of these results together, we can simplify our simulation model (Section 2.3) to formulate a minimal mechanistic model of MPB dynamics. Because the large-tree preference and tree size-dependent fecundity are non-essential (for excitability in MPB, specifically), we drop explicit age structure and only track the density of all susceptible trees (i.e., trees with DBH $>$ 8 inches). All susceptible trees are equally preferred for mass attacks, and all killed trees give rise to the same number of beetles. For simplicity, stochasticity is eliminated. The minimal mechanistic model can be written as
@@
$\underbrace{S_{y+1}}_{\mathclap{\text{Susceptible trees per acre}}}=\overbrace{g(K^{*}-S_{y})}^{\mathclap{\text{Recruitment of susceptible trees}}}-\underbrace{A_{y}\times(1-e^{\frac{-2.3A_{y}}{a}})\times\overbrace{\gamma^{*}}^{\mathclap{\text{Trees killed per succ. beetle attack}}}}_{\mathclap{\text{Trees killed by MPB}}}.$ (15)
@@
Here, $A_{y}$ is the female attack density in the $y^{\text{th}}$ year, $S_{y}$ is the density of susceptible trees, $S_{y}^{\prime}$ is the density of susceptible trees remaining after beetle attack (i.e., $S_{y}^{\prime}=S_{y}-A_{y}\times(1-e^{\frac{-2.3A_{y}}{a}})$), the parameter $\zeta^{*}$ is the average emerging females per killed tree, $\gamma^{*}$ is the average number of killed trees per successful attack, $K^{*}$ is redefined as the carrying capacity of the susceptible trees, $\epsilon$ is the now-explicit number of endemic female beetles, $g$ modulates the rate of recruitment into the class of susceptible trees, and all other parameters have their original meanings (see Table LABEL:tab:model_notation). When simulating the model, a floor at $S=0$ may be necessary to avoid negative trees.
@@
The minimal model has several purposes. 1) It is descriptive, compactly representing the density-dependent dynamics of MPB and lodgepole pine. 2) It demonstrates that there are two stable equilibria: an endemic equilibrium with few beetles, and an epidemic equilibrium with many beetles. This bistability is typical of systems with a strong Allee effect (Taylor,, 2012), so it is no surprise that the endemic equilibrium disappears when the Allee effect is turned off (Fig. 10C) 3) It shows the MPB/lodgepole system is an instance of an excitable dynamical system, where perturbations can lead to long excursions in phase space, before returning to a resting state (Fig. 10A). Typically, excitable systems have a single stable equilibrium (Strogatz,, 1994, pg. 118), whereas our system has two. However, the endemic equilibrium is globally-attracting as a practical matter, since the epidemic equilibrium has a minuscule basin of attraction (Fig. E.7). Previous papers proposed two stable equilibria for irruptive bark beetles (e.g., Berryman,, 1979; Mawby et al.,, 1989; Brush and Lewis,, 2024). Our model corroborates this while adding crucial nuance: while it is possible to temporarily attain beetle densities at (or above) epidemic equilibrium, it is not possible to stay at the epidemic equilibrium.
@@
Figure 10: Phase portraits of the minimal mechanistic model (Eq. 14 & Eq. 15), across various simulation scenarios Panel A: Stand-dependent dispersal, an Allee effect, and time-scale separation leads to excitable dynamics. Nullclines show curves where $\log(A_{y+1})=\log(A_{y})$ and $S_{y+1}=S_{y}$, and arrow directions are $\log(A_{y+1})-\log(A_{y})$ and $S_{y+1}-S_{y}$. Most model parameters were set to their respective posterior means (see Table 2). The exceptions are $\gamma^{*}=1/360$ and $\zeta^{*}=1300$, which are approximately the respective means (across trees) of attacks and emergences in the Klein et al., dataset. The carrying capacity’s value, $K^{*}=130$, is approximately the density of trees with $\text{DBH}\geq 8\text{in}$ in the Klein et al., dataset. The recruitment parameter’s value, $g=0.036$, was obtained by solving the differential equation $dS(t)/dt=g(K-S(t))$ with initial condition $S(0)=0$, then solving for the $g$ where 90% of carrying capacity is attained in $t=50$ years; lodgepole pine attains an 8-inch DBH and becomes susceptible after approximately 50 years.
@@
Through the development and analysis of our age-structured simulation model, we have shed light on the life history and biological processes underlying outbreaks, and challenged established ideas in the MPB modelling literature. A comparative analysis of bark beetles reveals that MPB’s life history traits contribute to its uniquely destructive nature. However, to accurately predict total damage, it’s necessary to consider host-tree biomass (Fig. 9) and idiosyncratic factors, such as fungal symbionts and interspecific competition. Our findings are succinctly summarized in a minimal mechanistic model (Eq. 14 & Eq. 15), which establishes the MPB/lodgepole system as the first empirical examples of an excitable dynamical system in ecology. This work underscores the importance of detailed case studies and looking beyond traditional predator-prey models to understand complex population dynamics.